# Decision Theoretic Foundations for Experiments Evaluating Human Decisions

### v5

@@ Abstract.
Decision-making with information displays is a key focus of research in areas like human-AI collaboration and data visualization. However, what constitutes a decision problem, and what is required for an experiment to conclude that decisions are flawed, remain open to speculation. We present a widely applicable definition of a decision problem synthesized from statistical decision theory and information economics. We claim that to attribute loss in human performance to bias, an experiment must provide the information that a rational agent would need to identify the normative decision. We evaluate whether recent empirical research on AI-assisted decisions achieves this. We find that only 10 (26%) of 39 studies that claim to identify biased behavior present participants with sufficient information to make this claim in at least one treatment condition. We motivate the value of studying well-defined decision problems by describing a characterization of performance losses they allow us to conceive.
@@ 1. Introduction
Identifying effective, informative interfaces to aid decision-making is a common goal in research in human-centered artificial intelligence (HCAI), visualization, and related areas. For example, visualization researchers emphasize assisting decision-making as a vital goal of data visualization (Dimara and Stasko, 2021). Similarly, in human-centered AI, a proliferation of empirical studies on human decision behavior have been deemed “necessary to evaluate the effectiveness of AI technologies in assisting decision making, but also to form a foundational understanding of how people interact with AI to make decisions” (Lai et al., 2023). However, research areas in HCI and related fields that rely heavily on decision-making lack a consensus on what a decision is, and how decisions are a distinct form of human behavior compared to other tasks. Human decisions from prediction and other information displays are often evaluated in studies with the goal of identifying forms of bias: ways in which study participants’ behavior deviates from expectations of good decision-making. But reviewing the range of tasks that get filed under decision-making, it is difficult to ascertain what common task structure makes these studies about decision-making, much less what properties a study needs to have in order for it to be capable of evaluating human decisions, rather than simply describing them.
@@
As examples of the inconsistency and underdefinition of good decision-making in the HCI literature, colloquial definitions describe a decision task as a choice between alternatives, or a choice where the stakes are high (such as deciding which of multiple hurricane forecasts is more accurate (Padilla et al., 2015), or whether to screen child welfare cases (Zytek et al., 2021)). Neither isolates a clear set of assumptions that must hold for a judgment to qualify as a decision. Other accounts cast decision-making as a higher-level task akin to sensemaking (Dimara and Stasko, 2021). For some, the point of studying human decisions is to compare them to a definition of idealized behavior. Sometimes researchers define this standard in a way that is common across studies (e.g., a better forecast comes closer to the true probability of the event (Green and Chen, 2019; Fogliato et al., 2021)). Other times, however, different definitions of the same concept are used: e.g., appropriate reliance as agreeing with an AI when it is correct and disagreeing when it’s not (Bansal et al., 2021; Buçinca et al., 2021), versus providing a higher subjective rating of trust when the model is correct (Bussone et al., 2015), versus choosing the better of the human versus the AI prediction to bet on for a future task (Dietvorst et al., 2015). Sometimes the point is to study seemingly “subjective decisions,” like how people react to music recommendations (Kulesza et al., 2012), or decide whether a house seems to match its valuation (Prabhudesai et al., 2023), where it can seem unclear whether normative interpretations of task performance can be applied. Not only is a more coherent definition of decision needed, as recent survey papers on decision making in visualization and human-centered AI recommend (Dimara and Stasko, 2021; Lai et al., 2021), but such a definition must clearly define requirements for evaluating decisions by identifying forms of bias or performance loss in human decision-making.
This paper contributes a definition of a decision problem and associated guidelines for practice for the purposes of evaluating human decisions from information displays.
We show how a widely applicable framework for studying decision-making can be snthesized from foundations of statistical decision theory and expected utility theory (Savage, 1972; Steele and Stefánsson, 2015; Von Neumann and Morgenstern, 2007), which rigorously motivate what it means to make good decisions under uncertainty, and information economics, which formalizes information structures that capture the relevant information in a set of signals for a decision problem (Bergemann and Morris, 2019). We argue that for a decision problem to be well-defined, it must be possible to rigorously define a normative (best response) decision to the information provided to the study participants. We motivate this criterion by showing how performance loss can only be conceived of when a study participant could in theory identify how to optimize their behavior from the information they are provided. To assess how close recent practice comes to achieving this ideal, we apply our definition to 46 recent studies on AI-assisted decision-making, finding that while 39 of the studies draw conclusions about shortcomings in human decisions with prediction displays, only 10 (26%) of these provided participants a well-defined decision problem in at least one treatment condition.
@@
This paper is organized as follows. First, we define a decision problem, normative decision, and sources of performance loss given human decision-makers. Next, we discuss considerations of scope and important study design decisions in applying this framework in practice. We then present two example study designs in detail, and describe how they could be improved to align with the framework we present. We apply the framework to a sample of studies on AI-assisted decisions, illustrating the gap between the ideal it presents and current practice. We conclude with recommendations for experimenters and a discussion of common misunderstandings of the approach we present.
@@ 2. Defining a Decision Problem
We define a decision problem and the normative decision given such a problem and enumerate possible sources of loss.
@@ 2.1. Decision Problem
Under this definition, the formalization of the joint distribution enables a definition of normative behavior. Summarizing information about , (as well as ) provides a basis for the experimenter to establish expectations under optimal use of the information.
@@ 2.2. Normative Behavior
Given a decision problem defined as above, we calculate the normative (“optimal”) decision by making assumptions about what it means for an agent to have coherent preferences and to optimally decide between actions under uncertainty about the outcome. Consequently, we can interpret an experiment participant’s performance as an attempt at achieving this standard, and identify sources of error (loss) in performance.
@@ 2.2.1. Rational Belief Formation
Specifically, to calculate the optimal decision for a decision task we first define what the agent perceives the probability distribution over the state to be prior to obtaining information from : or as we describe above. Whenever the signaling policy does not reveal directly via the signal, but does inform on , we assume that after seeing the signal, the agent updates their prior belief to a posterior belief over signals and states, using Bayes rule based on what they have learned about the data-generating model :
where is a normalizing factor.
The maximally informative posterior beliefs that any agent can arrive at given the decision problem is:
Note that the above definition differs from the posterior definition in Equation 1 only by explicitly assuming that the agent knows . It is sometimes unreasonable to expect that can be communicated to any agent, rational or not, without incurring some error. For example, in many studies that involve machine learning predictions based on many features and possibly many states (e.g., multi-label classification), communicating the joint distribution over signals and states without loss of information is impractical. Under such conditions, we expect the rational agent to arrive at posterior beliefs defined by:
@@ 2.2.2. Defining the Normative Action
Given the posterior beliefs as defined in Equation 2 or 3, we use Equation 5 to identify the action that a fully rational agent would choose in order to maximize their expected utility under . This is the normative decision.
To do so, we will first assume that the agent’s preferences can be summarized by the scoring rule . We assume that she will choose the action that maximizes her expected utility (score):
@@
We can define the optimal action as that which maximizes the agent’s expected utility:
@@ 2.3. Example
The data-generating model is the joint distribution over the state (delay or not) and signals comprised of flight feature values and AI prediction information. Assume that participants are presented with a signal in the form of a set of feature values describing a flight (origin, destination, time, carrier, prior delays in last 30 days, etc.) from which they must decide on an action. Imagine that the signal also includes the prediction from a model that participants are told was trained on a historical dataset of labeled examples.
For a participant to in theory be able to form posterior beliefs about the probability that a flight is delayed after viewing such signals, they must have prior information about the data-generating model. The joint distribution assigns a probability, for each possible value of ({0, 1}), to each possible signal (a combination of feature values and model prediction) used as a trial in the experiment. Rather than attempting to communicate the exact prior to participants, imagine that before beginning the study trials, participants are provided with a set of labeled instances, which defines . If we did not show the agent any feature values at all (nor model predictions), we would expect them to make their choice based on the probability that a randomly drawn flight in the study will be delayed, , representing the projection of to the state.
@@ 2.4. Defining Sources of Loss
Agent obtains prior
Agent processes signal
Agent updates beliefs
Agent takes optimal action
The four steps above identify four sources of loss representing possible reasons why observed human performance in an experiment might not match the normative standard. A final source of loss that must be considered when dealing with human (as opposed to hypothetical rational) agents is misunderstanding of the decision problem.
@@
3. Applying the Normative Decision Framework in Practice
We outline the scope of the above definitions in terms of what kinds of human subjects studies they apply to, and describe the primary implications for researchers seeking to evaluate human decisions. We discuss considerations for the design of the signal and scoring rule.
3.1. Scope
Our definition of a decision problem and corresponding standard for optimal behavior is intended for controlled evaluation of human behavior, a.k.a. normative decision research. Behavioral data (which may be produced by humans or by simulation) is collected under controlled conditions, with the goal of learning about behavior induced by the provision of information. Such studies are frequently used to describe the quality of human performance in some situation (e.g., how well people make decisions from displays in strategic settings (Zhang et al., 2023)), to rank different assistive elements by human performance (e.g, different visualizations (Correll et al., 2018) or AI explanation strategies (Liu et al., 2021)), or to test a hypothesis about how humans make decisions or what will help them do better (e.g., cognitive forcing functions will improve AI-assisted decisions (Buçinca et al., 2021)). Critically, such evaluative studies require that ground truth can be identified for any hypothetical states of the world against which study participants’ decisions are evaluated.
It is important to recognize that even if the researchers do not think about the elements of their experiment in terms of the components identified above (i.e., statespace, data-generating model, signaling policy, action space, scoring rule), the design of their study still induces a decision problem whenever two simple conditions hold (Wu et al., 2023). First, there must be some uncertainty about the state (i.e., there is more than one data scenario used to generate stimuli)111A single trial study where all participants view the same signal would not qualify., and second, the researchers believe it is possible to provide a scoring rule that is consistent across participants. This requirement eliminates, for example, studies of emotional responses.
3.2. Implications for communicating decision tasks to study participants
@@
The primary motivations for using the above framework are epistemological in nature: they concern what knowledge can be gained from the results of a study. Our primary claim in this work is that to evaluate human decision-making, an experiment must provide the participant with enough information to in principle identify the normative decision, which is used to judge their behavior. In other words, does the experiment give the participant enough information to align their understanding of the decision problem with its normative interpretation? Researchers must communicate to participants the action space, state space, scoring rule, and sufficient information about the data-generating model for participants to in theory arrive at the posterior beliefs against which they are evaluated.
@@
When the experiment provides the participant sufficient information, we can characterize forms of bias as loss in decision performance according to the four possible sources described in Section 2.4. When an experiment does not provide sufficient information for the participant to solve the task in principle, there is the potential for multiplicity concerning the data-generating model. Such multiplicity has implications for interpreting the study results, as we demonstrate through examples in Section 2.3.
3.3. Designing signals and feedback
@@
There is often more than one way for a study to provide sufficient information for a rational decision-maker to identify the normative decision. Experiment designs can vary in the amount of information they provide participants about or how they provide this information, provided enough information is conveyed for a rational agent to identify the best response.
@@
For example, for many experiments, it is critical to convey or an approximation of as the prior in order for participants to be able to update these beliefs upon viewing signals. However, for some experiments, it is instead possible for the signal to directly inform of the posterior probability of the state. For example, in a visualization study that asks users to decide whether or not they want to purchase a piece of special equipment expected to improve their score in a game (e.g., (Hofman et al., 2020; Kale et al., 2020)), a visualization of the player’s expected score distribution with and without the equipment (from which the realized outcome is drawn depending on their choice) conveys the posterior probability of the state directly.
Some experiment designs support learning about by providing participants feedback informing them of the realized state after each decision trial. When feedback is provided, we would expect the Bayesian rational decision-maker to update their beliefs according to the information provided on the joint distribution between signals and states. Imagine that participants in the flight booking study see feedback (i.e., are told the realized state, whether or not the flight was delayed) after each decision they make. Assuming the signal is not revealing of the posterior probability of the state (delayed or not), we would naturally expect the participants to learn about from this feedback as they complete trials, affecting their posterior beliefs (Equations 1 and 2). A challenge in using this approach without providing other information about (i.e., through signals and the prior) is that how much a participant can learn from any single trial is quite limited. Hence, such designs tend to require many trials.
3.4. Choice of Scoring Rule
As previously mentioned, a critical yet often poorly understood component of a decision problem is a scoring rule : that maps the action and state to a quality or payoff. A scoring rule has two possible uses: to incentivize participants and to evaluate the quality of the decision responses.
Whether the researcher uses the same scoring rule for both incentives and evaluation versus two different rules has important implications for interpreting the results. For the flight booking example, if the researchers analyzed the collected responses using the above rule (i.e., reporting performance in payoff space) in addition to communicating it to participants for incentivization, then the utility associated with performance is defined by a single scoring rule, and the interpretation of a “good” decision is consistent across incentivization and evaluation, which is clearly beneficial for interpreting the results.
If we imagine instead that the researchers decided to consider errors on a more equal footing in analyzing their data, for example reporting the results in terms of accuracy, where each type of error is associated with the same cost, the meaning of a given action differs between incentivization and evaluation. A scoring rule induces a meaning on actions by defining a range of beliefs for which an action is optimal. Consequently, interpreting results is difficult whenever two different rules are used. One rule will by definition “misinterpret” the actions of the other by implying a different normative standard.
One way around this problem when using two different rules is to make sure that the rule used for incentivization has certain properties. Specifically, among scoring rules for beliefs, we can distinguish those that are proper scoring rules: scoring rules with for which the optimal action is to predict the true distribution, i.e., . When a proper scoring rule is used, the researcher can assume that an agent who understands the rule properly will report their true beliefs, as it is not possible to score higher by reporting some alternative beliefs. These beliefs can then be plugged into a different scoring rule without the risk of misinterpretation.
An important implication of the proper scoring rule distinction for belief reports is that in general we should not expect an agent to truthfully report their beliefs when it is not in their best interest to do so. Consequently, when non-proper rules are used for belief tasks, we cannot interpret a study participant who does not report the appropriate beliefs after examining the signal as demonstrating bias. Such “deviation” from what the researcher might expect (honest reporting) is not technically a loss, it should be the expected behavior under that rule.
Applying the construction in Equation 7 to a binary decision gives a threshold rule where for each state, all reported probabilities below a threshold receive one score and those above receive another score. Since the payoff depends only weakly on the belief, such scoring rules may be less sensitive to the agents’ understanding of the signal. For this reason, including a decision problem that directly also elicits beliefs using a strictly proper-scoring rule is a good idea whenever eliciting responses using an action space that is not over beliefs, as it enables plugging the elicited beliefs into other scoring rules.222See, e.g., Wu et al. (2023)’s analysis of (Kale et al., 2020).
@@ 3.4. Choice of Scoring Rule
The choice of possible scoring rules depends on the nature of the action space. In general, a coarser action space (i.e., with fewer choices available to participants) will produce less informative responses for the purpose of resolving observed performance relative to a normative standard. If, for example, the researcher designed the flight booking study to also elicit the agent’s prediction of the probability of delay, each trial would provide more information, and they could employ a proper scoring rule such as squared loss (i.e., Brier or quadratic rule) or log score to score the accuracy of the probabilistic beliefs. Different proper scoring rules imply different expectations about the set of possible decisions (Levinstein, 2017). See Li et al. (2022) for a discussion of scoring rules that have a high information value for the rational agent.
4. Examples

### v6

@@ Abstract.
Decision-making with information displays is a key focus of research in areas like human-AI collaboration and data visualization. However, what constitutes a decision problem, and what is required for an experiment to conclude that decisions are flawed, remain imprecise. We present a widely applicable definition of a decision problem synthesized from statistical decision theory and information economics. We claim that to attribute loss in human performance to bias, an experiment must provide the information that a rational agent would need to identify the normative decision. We evaluate whether recent empirical research on AI-assisted decisions achieves this standard. We find that only 10 (26%) of 39 studies that claim to identify biased behavior presented participants with sufficient information to make this claim in at least one treatment condition. We motivate the value of studying well-defined decision problems by describing a characterization of performance losses they allow to be conceived.
@@ 1. Introduction
Identifying effective information displays to support data-driven inference and decision-making is a common goal across a number of domains, including research related to human-computer interaction (HCI). For example, visualization researchers emphasize assisting decision-making as a vital goal of data visualization (Dimara and Stasko, 2021). Similarly, in human-centered artificial intelligence (AI), a growing body of empirical studies on human decision behavior has been called “necessary to evaluate the effectiveness of AI technologies in assisting decision making, but also to form a foundational understanding of how people interact with AI to make decisions” (Lai et al., 2023). Human decisions from predictions and other information displays are often evaluated in studies with the goal of identifying forms of bias: ways in which study participants’ behavior deviates from expectations of good decision-making. However, within HCI and related fields there is little consensus on what a decision is, and how decisions are a distinct form of human behavior compared to other tasks.
@@
For example, colloquial definitions describe a decision task as a choice between alternatives, or a choice where the stakes are high (such as deciding which of multiple hurricane forecasts is more accurate (Padilla et al., 2015), or whether to screen child welfare cases (Zytek et al., 2021)). Neither isolates a clear set of assumptions that must hold for a judgment to qualify as a decision. Other accounts cast decision-making as a higher-level task akin to sensemaking (Dimara and Stasko, 2021). For some, the point of studying human decisions is to compare them to a definition of idealized behavior. Sometimes researchers define this standard in a way that is common across studies (e.g., a better forecast comes closer to the true probability of the event (Green and Chen, 2019; Fogliato et al., 2021)). Other times, however, different definitions of the same concept are used: e.g., appropriate reliance as agreeing with an AI when it is correct and disagreeing when it’s not (Bansal et al., 2021; Buçinca et al., 2021), versus providing a higher subjective rating of trust when the model is correct (Bussone et al., 2015), versus choosing the better of the human versus the AI prediction to bet on for a future task (Dietvorst et al., 2015). Sometimes the point is to study seemingly “subjective decisions,” like how people react to music recommendations (Kulesza et al., 2012), or decide whether a house seems to match its valuation (Prabhudesai et al., 2023), where it can seem unclear whether normative interpretations of task performance can be applied.
Not only is a more coherent definition of decision needed, as recent survey papers recommend (Dimara and Stasko, 2021; Lai et al., 2021), but such a definition must clearly define requirements for evaluating decisions by identifying forms of bias or performance loss in human decision-making. In particular, we are concerned with the minimum requirements for a controlled study of human behavior to support claims that the human decision-making is flawed or non-ideal in some way. While it is generally recognized that such evaluative studies require ground truth outcomes to be identifiable for decision instances provided to participants, methodological work on experimental design in HCI and beyond lacks much formal guidance on what constitutes a valid research design for drawing normative conclusions about human decisions.
This paper contributes a definition of a decision problem and associated guidelines for experiment design that apply to any study intended to evaluate human decision-making under the provision of information. We show how a widely applicable framework for studying decision-making can be synthesized from several bodies of work outside HCI. First, foundations of statistical decision theory and expected utility theory (Savage, 1972; Steele and Stefánsson, 2015; Von Neumann and Morgenstern, 2007) rigorously define what it means to make good decisions under uncertainty. Second, information economics formalizes information structures that capture the relevant information in a set of signals for a decision problem (Bergemann and Morris, 2019). We argue that for a decision problem to be well-defined, it must be possible to define a normative (best response) decision to the information provided to the study participants. In other words, whether the results of a study can be used to argue that humans are biased depends on the answer to the question: Could a rational agent who perfectly perceived the information have used it to identify a best response? Importantly, this best response is calculated relative to the information available to the participant. This makes it distinct from scoring responses directly against the ground truth outcome. As an example of the difference, imagine a study where participants are shown weather forecasts, and asked to make decisions about whether to bring an umbrella or not. Rather than concluding that a participant made the wrong decision to bring the umbrella because a simulated outcome was no rain, they should be scored relative to whether the choice they made maximized their expected utility according to the scoring function the researcher uses to tabulate results. Additionally, this scoring function should have been communicated to them if the researchers want to draw conclusions about how good their decisions were.
We motivate this criterion by showing how performance loss can only be conceived of when a study participant could in theory identify how to optimize their behavior from the information they are provided. If, for example, the researchers had not communicated to participants of the weather forecast study sufficient information about how their responses would be judged, then we should expect different participants to make different guesses about what their goals should be. Consequently, their responses are incomparable, and it is left ambiguous what decision problem they thought they were solving. On the other hand, if they were provided sufficient information on the scoring process and provided forecast information, it is in theory possible to distinguish between a limited set of errors that explain observed performance loss, such as not extracting all of the decision-relevant information from the display, or not being able to identify the best response given the information they did extract.
To assess the extent to which recent empirical evaluations in HCI achieve this ideal, we apply our definition to 46 recent studies on AI-assisted decision-making, finding that while 39 of the studies draw conclusions about shortcomings in human decisions with prediction displays, only 10 (26%) of these studies provided participants with a well-defined decision problem in at least one treatment condition.
The benefits of adopting a more rigorous standard based in established theoretical foundations for decision-making (Savage, 1972; Bergemann and Morris, 2019) for empirical decision research in HCI, human-centered AI/ML, visualization, and related fields are twofold. First, it stands to improve the validity of interpretations of individual studies. The variation induced when a decision problem is not clearly reported nor communicated to participants confounds our ability to interpret the results, in that it cannot be separated from variation induced by true differences in how people do a task under different conditions. Adopting a standard where the normative decision arises from a well-defined process puts researchers in a better position to point to possible sources of performance loss, and raises awareness of the limitations on what can be concluded from study results when a decision problem is under-communicated. Second, it paves the way for the development of theory. Only when a decision problem is well-defined can human responses to that problem be interpreted relative to the results of other studies (e.g., that explore variations on the problem). This provides HCI researchers an ability to make meaningful contrasts to drive theory development. Otherwise, any apparent differences in the quality of behavior may reflect noise induced by presenting participants with an underspecified problem.
@@
The paper is organized as follows. First, we define a decision problem, normative decision, and sources of performance loss for human decision-makers. Next, we provide study design recommendations for applying this framework in practice. We present two example study designs in detail, and describe how they could be improved to align with the framework we present. We apply the framework to a sample of studies on AI-assisted decisions, illustrating the gap between the ideal it presents and current practice. We conclude with a discussion of limitations and future work.
@@ 2. Defining a Decision Problem
We define a decision problem and the normative decision given such a problem and enumerate possible sources of loss. Our definitions are intended for controlled evaluation of human behavior, a.k.a. normative decision research. Behavioral data (which may be produced by humans or by simulation) is collected under controlled conditions, with the goal of learning about behavior induced by the provision of information. Such studies are frequently used to describe the quality of human performance in some situation (e.g., how well people make decisions from displays in strategic settings (Zhang et al., 2023)), to rank different assistive elements by human performance (e.g, different visualizations (Correll et al., 2018) or AI explanation strategies (Liu et al., 2021)), or to test a hypothesis about how humans make decisions or what will help them do better (e.g., cognitive forcing functions will improve AI-assisted decisions (Buçinca et al., 2021)). Critically, such evaluative studies require that ground truth can be identified for any hypothetical states of the world against which study participants’ decisions are evaluated.
@@ 2.1. Decision Problem
Under this definition, the formalization of the joint distribution enables a definition of normative behavior. Summarizing information about , (as well as ) and communicating it to the participant provides a basis for the experimenter to establish expectations under optimal use of the information.
@@ 2.2. Normative Behavior
Given a decision problem defined as above, we calculate the normative (“optimal”) decision by assuming that the agent has coherent preferences under a scoring rule (dictated by a small set of axioms (Von Neumann and Morgenstern, 2007)) and uses them to decide between actions under uncertainty about the outcome. Consequently, we can interpret an experiment participant’s performance as an attempt at achieving this standard, and, if sub-optimal, identify sources of error (loss) in performance.
@@ 2.2.1. Rational Belief Formation
Specifically, to calculate the optimal decision for a decision task we first define what the agent perceives the probability distribution over the state to be prior to obtaining information from : or as we describe above. Whenever the signaling policy does not reveal directly via the signal, but does inform on , we assume that after seeing the signal, the agent updates their prior belief to a posterior belief over signals and states, using Bayes rule based on what they know about the data-generating model . When the signaling policy does reveal directly via the signal (i.e., and ), this update is trivial. The maximally informative posterior beliefs that any agent can arrive at given the decision problem is:
It is sometimes unreasonable to expect that can be communicated to any agent, rational or not, without incurring some error. For example, in many studies that involve machine learning predictions based on many features and possibly many states (e.g., multi-label classification), communicating the joint distribution over signals and states without loss of information is impractical. Under such conditions, we expect the rational agent to arrive at posterior beliefs defined by:
@@ 2.2.2. Defining the Normative Action
Given the posterior beliefs as defined in Equation 1 or 2, we use Equation 4 to identify the action that a fully rational agent would choose in order to maximize their expected utility under . This is the normative decision.
We assume that the agent’s goal is to maximize their expected score under the scoring rule :
@@
The optimal action is then the action that maximizes the agent’s expected score:
2.2.3. Eliciting and Scoring Beliefs
A scoring rule induces a “meaning” on actions by demarcating a range of beliefs for which an action is optimal. Consequently, when different rules are applied to incentivize participants versus to evaluate their responses, the meaning of a given action differs. One rule will by definition “misinterpret” the actions of the other by implying a different normative standard. This misinterpretation implies that the scoring rule should be consistent across the two purposes of incentivizing participants and evaluating their responses.
An exception to this requirement occurs when beliefs are elicited instead of decisions, and the rule used to incentivize belief reports has certain properties. Among scoring rules for beliefs, we can distinguish those that are proper scoring rules: scoring rules with for which the optimal action is to predict the true distribution, i.e., . Under a proper scoring rule, it is not possible to score higher by reporting some alternative beliefs. These beliefs can then be plugged into a different proper scoring rule without the risk of misinterpretation.
For example, applying the construction in Equation 5 to a binary decision gives a threshold rule where for each state, all reported probabilities below a threshold receive one score and those above receive another score.
2.3. Scope of the definitions
The above definitions are applicable to any studies that aim to produce normative conclusions about human behavior. In such studies, behavioral data (produced by humans or simulation) are collected with the goal of learning about how behavior is induced by the provision of information. Such studies are frequently used to describe the quality of human performance in some situation (e.g., how well people make decisions from displays in strategic settings (Zhang et al., 2023)), to rank assistive elements by human performance (e.g., different visualizations (Correll et al., 2018), AI explanation strategies (Liu et al., 2021), or prompts (Buçinca et al., 2021; hullman2017imagining)), or to test a hypothesis about how humans make decisions (Kale et al., 2020).
It is important to recognize that even if the researchers do not think about the elements of their experiment in terms of the components identified above (i.e., statespace, data-generating model, signaling policy, action space, scoring rule), the design of their study still induces a decision problem whenever two simple conditions hold (Wu et al., 2023). First, there must be some uncertainty about the state (i.e., there is more than one data scenario used to generate stimuli)111A single trial study where all participants view the exact same signal would not qualify., and second, it must be possible to score decision quality consistently across participants. This requirement eliminates, for example, studies of emotional responses.
@@ 2.4. Example
The data-generating model is the joint distribution over the state (delay or not) and signals comprised of flight feature values and AI prediction information. Assume that participants are presented with a set of feature values describing a flight (origin, destination, time, carrier, prior delays in last 30 days, etc.) from which they must decide on an action. Imagine that the signal also includes the prediction from a model that participants are told was trained on a historical dataset of labeled examples.
For a participant to in theory be able to form posterior beliefs about the probability that a flight is delayed after viewing such signals, they must have prior information about the data-generating model. The joint distribution assigns a probability, for each possible value of ({0,1}), to each possible signal (a combination of feature values and model prediction) used as a trial in the experiment. Rather than attempting to communicate the exact prior to participants, imagine that before beginning the study trials, participants are provided with a finite set of labeled instances , which defines a joint distribution over . Note that if we did not show the agent any feature values at all (nor model predictions), we would expect them to make their choice based on the prior probability that a randomly drawn flight in the study will be delayed, .
@@ 2.5. Defining Sources of Loss
Consider the process implied by definitions in Sections 2.1-2.2, and how each step might result in lost performance relative to a rational agent:
Agent obtains prior prior loss
Agent processes signal receiver loss
Agent updates beliefs updating loss
Agent takes optimal action optimization loss
For these four sources of loss to explain observed performance losses from study participants, we assume the participants understand the decision problem described to them. However, it always possible that they do not, implying a fifth possible source of loss.
@@
3. Recommendations
@@
Our primary argument in this work is that to evaluate human decision-making, an experiment must provide the participant with enough information to in principle identify the normative decision, which is used to judge their behavior. In other words, does the experiment give the participant enough information to align their understanding of the decision problem with its normative interpretation?
The primary motivations for using the framework are epistemological in nature: they concern what knowledge can be gained from the results of a study. A lack of consensus around how to define a decision problem and the minimum required components for normative analysis leads to research landscape in which researchers can easily be tempted to overinterpret the results of their studies. Readers may not recognize the overinterpretations, leading to a research literature based on invalid claims about human biases.
@@
In contrast, when the experiment provides the participant sufficient information, we can characterize forms of bias as loss in decision performance according to the four possible sources described in Section 2.5. When an experiment does not provide sufficient information for the participant to solve the task in principle, there is the potential for multiplicity concerning the data-generating model, as we cannot assume that the participant understood the data-generating model as the researcher intended. Consequently, researchers wishing to design a normative decision experiment should communicate all necessary components of the decision problem both to study participants and in publications. Specifically, we recommend that researchers:
Clearly convey action and state spaces and a scoring rule. The finite action space and statespace should be clearly communicated to participants, and the action space reflected in the instruments used to collect participant responses. Because the optimal action can only be determined relative to a scoring rule : that maps the action and state to a quality or payoff, participants should be given a scoring rule to reduce heterogeneity in their choice of action. This is necessary for interpretation even if the experimenter feels comfortable assuming that the participants will try their best and does not use the scores in assigning monetary rewards.
@@
Provide prior and/or sufficient information about the data-generating model to calculate posterior. As discussed above, sufficient information about the data-generating model must be provided to arrive at the Bayesian posterior distribution. In designing experiments, researchers should recognize that there is often more than one way for a study to provide sufficient information for a rational decision-maker to identify the normative decision. However, the fact that experiment designs can vary in the amount of information they provide participants about or how they provide this information does not mean that anything goes when it comes to what information is provided. Regardless of the specifics, sufficient information for a rational agent to identify the best response must be provided to interpret the results.
@@
For many experiments, it is critical to convey or an approximation of as the prior in order for participants to be able to update these beliefs upon viewing signals. However, for some experiments, it is instead possible for the signal to directly inform of the posterior probability of the state. For example, in a visualization study that asks users to decide whether or not they want to purchase a piece of special equipment expected to improve their score in a game (e.g., (Hofman et al., 2020; Kale et al., 2020)), a visualization of the player’s expected score distribution with and without the equipment (from which the realized outcome is drawn depending on their choice) conveys the posterior probability of the state directly.
Whenever the signal on its own does not induce a posterior, the prior should be endowed, and sufficient information about should be provided to enable estimating and . Exemplar-based training procedures such as have been used to teach humans when to defer to an AI model can be employed for this purpose (mozannar2022teaching). More generally, using presentation formats that are more likely to be automatically processed (e.g., conveying distribution information via frequency animations (Goldstein and Rothschild, 2014; Hullman et al., 2015)) or validating participants’ priors through elicitation (e.g., (Kim et al., 2019, 2020)) can help researchers understand and minimize prior loss.
Use the same scoring rule for incentivization and evaluation, or elicit beliefs with a proper scoring rule. As described above, using different scoring rules for incentivizing participants versus evaluating their responses means that “good” behavior is defined differently across these two activities, confounding interpretation. Researchers should use the same rule to avoid this confound.
An exception to this rule applies when beliefs are elicited using a proper scoring rule but behavior is analyzed using a different, downstream rule. There are several benefits to eliciting beliefs using a proper scoring rule, in place of or in addition to eliciting decisions. First, when a proper scoring rule is used, the researcher can assume that an agent who understands the rule properly will report their true beliefs. This supports interpretation by eliminating ”deviation by design,” referring to how deviation from honest reporting is not technically a loss under a non-proper rule, it should be the expected behavior under that rule. In other words, in a non-proper scoring rule, we cannot interpret as biased the behavior of a study participant who does not report the appropriate beliefs after examining the signal. Eliciting beliefs using a strictly proper scoring rule enables plugging the elicited beliefs into other scoring rules.222See, e.g., Wu et al. (2023)’s analysis of (Kale et al., 2020).
@@
Second, belief elicitation through proper scoring rules can provide benefits in the form of more information captured in the elicited responses. In general, a coarser action space (i.e., with fewer choices available to participants) will produce less informative responses for the purpose of resolving observed performance relative to a normative standard. For example, if the researcher designed the flight booking study to also elicit the agent’s beliefs, in the form of their prediction of the probability of delay, each trial would provide more information, and they could employ a proper scoring rule such as squared loss (i.e., Brier or quadratic rule) or log score to score the accuracy of the probabilistic beliefs.
When eliciting beliefs, one should also acknowledge that different proper scoring rules imply different expectations about the set of possible decisions (Levinstein, 2017). See Li et al. (2022) for a discussion of scoring rules that have a high information value for the rational agent.
Compare to best attainable performance. We observed that even when researchers invest effort to design scoring rules to incentivize participants, they do not often compare the observed behavioral performance to the best attainable performance under the rule. Researchers should default to analyzing results in the space of the scoring rule used for incentives, and provide a rationale whenever they deviate.
Account for learning when providing feedback. Some experiment designs support learning about by providing participants feedback informing them of the realized state after each decision trial. When feedback is provided, we would expect the Bayesian rational decision-maker to update their beliefs according to the information they gain about the joint distribution between signals and states. Imagine that participants in the flight booking study see feedback (i.e., are told the realized state, whether or not the flight was delayed) after each decision they make. Assuming the signal is not revealing of the posterior probability of the state (delayed or not), we would naturally expect the participants to learn about from this feedback as they complete trials, affecting their posterior beliefs (Equation 2). A challenge in using this approach without providing other information about (i.e., through signals and the prior) is that how much a participant can learn from any single trial is quite limited. Hence, such designs tend to require many trials if the goal is to study behavior under relatively stable beliefs about the data-generating process.
Elicit measures to estimate specific sources of loss. Assuming the same decision problem has been conveyed to participants as the researchers analyze, it becomes possible to conceive of a set of losses to aid interpretation of results. How to estimate sources of loss in an unbiased way is a rich area for future work. In many cases, we expect separating losses to entail adding carefully constructed questions to the experimental procedure, and/or constructing benchmarks based in rational agent performance to separate sources of loss in observed results. For example, Wu et al. (2023) use the notion of a calibrated behavioral agent along with a rational agent benchmark and baseline (expected performance of a rational agent under only the prior) to separate optimization and receiver loss. An important area for future theoretical work is to establish under what conditions the losses we conceptualize can be estimated.
Conduct design analysis. Researchers doing normative decision research can benefit from conducting analyses prior to running the study that simulate the idealized actions of an agent to vet the experiment design (see, e.g., Wu et al. (2023)’s pre-study analyses for visualization studies and Guo et al. (2024)’s analysis of the potential for complementarity in AI-assisted decisions). Simulation-based design analysis forces the researcher to explicitly define the parameters of their decision problem, which can elucidate what information needs to be communicated to a participant in order to enable an optimal response. It also enables the researcher to test variations on their data-generating model and scoring rule, which can help the researcher realize under what conditions (1) the experiment can have acceptable statistical power and (2) participants can be paid fairly. Design analysis helps the researcher choose conditions (e.g., AI explanations) that provide informative signals for the decision task, narrowing down what kinds of research questions can be answered using a particular decision problem. For example, at a minimum, to be informative of how effectively people use signals, a study design must have sufficient statistical power to separate the rational agent performance with and without the signal.
4. Examples: Using the Framework to Redesign Confounded Studies